Raman system status

Raman system status

by Andrei Sapelkin -
Number of replies: 0

Dear All,

Just a quick update on the Raman system status. I spent a couple of days last week assessing the system health and took the opportunity to bring the 442 nm wavelength back into operation as well as to optimise 785 nm laser. The alignment was successful and both 442 and 785 nm laser is now available for users (for your reference the measured full laser power on sample using 20x objective was 20 mW for 442nm and 52 mW for 785 nm). The 633 nm laser wavelength  is still not available. 

However, it came to my attention that there may still be intermittent operation of some parts of the systems including laser shutter and sample enclosure/microscope door. These, together with 633 nm laser will be addressed during Renishaw engineer visit due to take place on 14th and 15th February.

Meanwhile, please report any problems you may have encountered while using the system using user form in the system page in QMPlus.

Best wishes

Dr Andrei Sapelkin